OpenStudy (anonymous):

graph each function f(x)=log1/3(3-x).....trying to figure this out... i know its going to shift to the left 3 spots and will have the vertical asymptote at x=-3...but i think what i have so far is wrong...any help??

OpenStudy (anonymous):

0

OpenStudy (dumbcow):

whenever "x" is neg, pull out the negative so you can identify reflection and it helps not confuse the shift \[\log (3-x) = \log[-(x-3)]\] shift 3 to right ... asymptote at x=3 reflection about x=3

OpenStudy (dumbcow):

oh i missed something....the base is less than 1 so it has vertical reflection about y=0 as well \[\log_{1/b} x = -\log_{b}x \]

OpenStudy (dumbcow):

to check answer use graphing software http://www.wolframalpha.com/input/?i=plot+log_%281%2F3%29+%283-x%29

OpenStudy (anonymous):

thanks for your time.

OpenStudy (dumbcow):

no problem, glad to help